Sol Summary – December 1st

Sol: 4
Summary Title: Prepping for Flight
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: EVA excursion in the local area around HAB working on drone flight preparation, EVA suit repair in the RAM module, and plantings at the GreenHab.
Look Ahead Plan: EVA to conduct first drone flights above ground paths of recent excursions (e.g. Marble Ritual).
Anomalies in work: None
Weather: Calm morning, with clouds arriving by the afternoon. As the sun set under overcast skies, winds increased quickly, reaching a Strong Breeze with gusts exceeding 30 mph.
Crew Physical Status: Crew in good shape. Some mild morning congestion was reported.
EVA: EVA led by the Technology Officer with support from the XO and Health and Safety Officer. 1 hour duration using Spirit and Perseverance rovers, staying within range of the HAB as work was done towards the initial drone flight.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

Commander Report – December 1st

Sol: 4
Summary Title: Space Suit Repair & EVA #4
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars with sunny skies, light high-altitude clouds and calm winds throughout the morning and into the afternoon. In the late afternoon, winds picked up and clouds began to move in. The crew enjoyed a breakfast of cereal, oatmeal, homemade sourdough bread, and espresso. After a morning planning meeting, HSO collected crew health stats. Technology Officer employed help from the two licensed private/drone pilots on the crew (CDR, HSO) to complete drone flight planning preparations for the afternoon EVA. After exercise and a light lunch, CDR, Technology Officer, and HSO embarked on EVA #4, while the XO led IVA Comms. The EVA plan included drone operations along Cow Dung Rd to Galileo Rd, but early technical challenges with the drone caused the crew to deviate from the scheduled EVA plan, remaining close to the Hab to troubleshoot. Due to this deviation, the crew was called back early to the Hab. During the EVA, Engineer and Green Hab Officer successfully completed repair operations on EVA suit 7 in the RAM. All crew members enjoyed assisting the Green Hab Officer in planting carrots and onions in the Green Hab in the afternoon. Before dinner, CDR, Technology Officer, and HSO replanned the drone flights for tomorrow’s requested morning EVA. The three ham radio operators on the crew (CDR, Green Hab Officer, and HSO) studied how to deploy the ham radio antenna for tomorrow’s requested afternoon EVA. The crew enjoyed a dinner of mac and cheese with homemade sourdough bread and peanut butter chocolate chip cookies.

Operations Report – December 1st

SOL: 4

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 201.8 (Start), 201.9 (End)
Beginning charge: 100%
Ending charge: 92%
Currently charging: No

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: Yes
Hours: 250.3 (Start), 250.3 (End)
Beginning charge: 100%
Ending charge: 90%
Currently charging: No

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 30 gallons used today, 110 gallons used in total thus far
Water (static tank): 440 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer (Ashley Kowalski) and GreenHab Officer (Matthew Eby) performed repairs on EVA suit #7. Ashley and Matt located the soldering iron, solder, and wire in the RAM and performed the repairs in that module. In order to remove the old solder, they used the soldering iron to melt it off. This enabled them to push new wires through the connection openings and then resolder new wires to that area. They were able to manually peel off any of the additional glue around the previous connections. After cleaning up the connection points, all wire connections were successfully reconnected to the old charging socket (we used EVA suit #8 as an example to follow). Please see attached photo displaying the repaired connections (please see photo 001 attached). However, even with all of the wire connections reconnected correctly, the fans were still not turning on. Therefore, Ashley and Matt tested all of the connections with the multimeter and noticed that even though the battery was reading ~13 V, the battery power was not transferring to the on/off switch. Upon further investigation, they realized that the fuse was blown (please see photo 002 attached). Thus, they located spare fuse parts in the EVA suit repair kit and put a new fuse in the system, which solved the problem. The fans turned on, the charging port was reconnected, and EVA suit #7 is now cleared for use.

One additional point that was noticed during the repair was that the lower connection on the on/off switch (the one that is circled in photo 003 attached; it is the one that connects the battery power to the fans) is a bit loose. Ashley and Matt attempted to tighten this connection, but it is still a bit loose. Why is this important? They believe that the looseness of this connection might have been the reason that the fuse blew. As you can see in the video uploaded to the Google drive link, when that connection is in one position (not pushed all the way down, but pushed slightly inward), it reads an incorrect value on the multimeter. When the connection IS pushed all the way down (pushed outward), the correct, expected value is read on the multimeter. Ashley and Matt recommend replacing this switch in the future as this connection may cause problems for future crews.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Today, the crew joined the GreenHab Officer in the greenhouse and helped with planting carrots and onions. Additional information can be found in the GreenHab Report.
WATER USE: ~23 gallons (~5 gallons today)
Heater: On
Supplemental light: Off
Harvest: N/A

Summary of ScienceDome operations: N/A
Dual split: No heat or AC used. All systems off.

Summary of RAM operations: A soldering iron, solder, wire, and a spare EVA suit fuse were used in the RAM to repair EVA suit #7. The heater was turned on while the Crew Engineer and GreenHab Officer were in the RAM performing the repairs. The heater was turned off and unplugged upon exiting the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The Crew Engineer provided Mission Support with the dimensions of the microwave grease filter. Also, when the wind picked up this afternoon before the crew was able to attach the additional security rope to the front airlock outer door, the door swung upon. The crew received permission to re-secure the door. Finally, we received the air filter for the furnace, but the Crew Engineer has not yet installed it.

Journalist Report – December 1st

Crew 269 Journalist Report
Date: 1 Dec 2022
Sol: 4
Author: Barbara Braun, Health and Safety Officer

The crew started their morning with a quick breakfast of oatmeal, rehydrated fruit, coffee/tea, and sourdough bread that our Health and Safety Officer, Barbara Braun, baked the night before. After morning health checks, Crew Engineer Ashley Kowalski and Green Hab Officer Matt Eby gathered equipment to repair the charging plug on one of the spacesuits, while Commander Kristine Ferrone, Tech Officer Trevor Jahn, and HSO Barbara Braun prepared to conduct an EVA to test out the crew’s imaging drone. Executive Officer Alli Taylor provided comms support to the EVA crew from the Hab.

The EVA crew had mapped out several locations for taking imagery, but encountered technical issues in the field with the drone’s memory card and the battery for the drone controller. The team was able to charge the battery and work around the issues with the memory card, but by then the time allotted for activities near the first drone site had passed. The team therefore checked on the mirrors experiment before returning to the Hab, and will re-attempt the EVA tomorrow.

Meanwhile, Ashley and Matt were able to successfully repair the spacesuit and complete other maintenance activities. The three ham radio operators on the crew (Kristine, Matt, and Barbara) looked at the ham radio equipment in the Hab in preparation for next week’s balloon launch and for possibly conducting some ham radio contacts. The Green Hab Officer, with the enthusiastic help of the rest of the crew, planted carrots and onions in the Green Hab for future crews to enjoy. Other crew members exercised using the compact exercise equipment in the Hab and had quick showers, before the whole crew settled down to a dinner of macaroni and cheese.

GreenHab Report – <date – December 1st>

GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 66-deg F at time of watering. Range prior 24 hours: 66-deg F – 87-deg F
Hours of supplemental light: 0 hours
Daily water usage for crops: Approximately 5 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 225 gallons (Based on a 250 gallon capacity, visually 90-95% full)
Time(s) of watering for crops: 6:30 PM
Changes to crops: Seeded an area of carrots and an area of onion (see Narrative below).
Narrative: Today 1 ½ rows of Black Nebula Carrots were planted in between the rows of tomato plants. In addition, 2 small rows of Scarlet Bandit Onions were planted in the bed formally used for Radish (stock of radish seeds exhausted from the seed bank).
Harvest: None
Support/supplies needed: None needed at this time.

EVA Report – December 1st

EVA # 4

Author: Barbara Braun, Health and Safety Officer
Purpose of EVA: Crew training with small robotic drone
Start time: 12:00 pm
End time: 1:20 pm
Narrative: Trevor, Kristine, and Barbara made their way out the airlock for the drone checkout EVA today at approximately 12:05 pm. With them they brought the crew’s small drone and a tote bag of support equipment. The trio first stowed the spare drone batteries and other equipment in the Spirit rover, and proceeded on foot to an area slightly to the north-northwest of the Hab where they set up the drone for flight. At this point the crew discovered that the drone controller battery was at 28%, and deeming this too low a state of charge to be able to complete the flight safely, the crew connected a battery to the controller to let it charge. While waiting for the controller to charge, the crew took the Spirit and Perseverance rovers to the corner of the road to the Hab and Cow Dung Road in order to pre-stage the rovers at the first drone landing location. The crew proceeded on foot to the location of the mirror sample experiment nearby to check on the integrity of the mirrors, and noted a buildup of salt on one of the mirrors which should provide good data to the experimenters when the mirror is retrieved for examination tomorrow. The crew then proceeded on foot back to the Hab area to check on the state of charge of the controller. The controller had sufficient charge to complete the first flight, so the crew began loading the first flight plan into the drone. The drone SD card then gave an error, but the crew loaded a backup copy of the flight plan and was preparing to launch the drone. At this point mission support called to inquire about the rovers parked at the intersection of the road to the Hab and Cow Dung Road. After consulting with mission support, the team realized that the EVA request did not include sufficient detail about the drone flight plans and the overall plan for the EVA, and the EVA was called off for the day. The crew stowed the drone and its equipment near the Hab airlock and proceeded on foot to retrieve the two rovers. The crew parked the two rovers, noted ending hours and state of charge, left them unplugged, and proceeded into the airlock at 1:15 pm with all equipment.

Destination: Hab, Marble Ritual / mirror deployment location

Coordinates (use UTM WGS 84): 519000, 4251000

Participants: Commander (Kristine Ferrone), HSO (Barbara Braun), Technology Officer (Trevor Jahn)

Road(s) and routes per MDRS Map: Road to the Hab, Cow Dung Rd (0110)

Mode of travel: Rovers and walking

Copyright © The Mars Society. All rights reserved. | Main Site